OpenStudy (anonymous):

find the equation in standard form of the parabola with vertex at (0,0) and focus at (0,-3)

OpenStudy (kc_kennylau):

So it goes down... I'm guessing -(x+3)^2? P.S. I'm not sure

OpenStudy (kc_kennylau):

ok i'm going for -3x^2

OpenStudy (kc_kennylau):

no it's neither lol

OpenStudy (anonymous):

the options are: y^2=12x x^2=-12y y^2=-12x x^2=12y

OpenStudy (kc_kennylau):

\[y=-\frac{1}{12}x^2\]

OpenStudy (kc_kennylau):

wait what

OpenStudy (kc_kennylau):

ok i'm not familiar in this kind of thing lol

OpenStudy (anonymous):

Thanks for helping anyway.